OpenStudy (anonymous):

At what POINTS is this function continuous..y=cosx/x...someone please explain to me how to solve these types of problems? is it whatever numbers make this an undefined problem?

OpenStudy (anonymous):

well x=/=0

OpenStudy (anonymous):

i would say limit x>0 \[(-\infty,0)\]\[(0,\infty)\]

OpenStudy (anonymous):

but im a nub so dont trust me much

OpenStudy (anonymous):

the book says all x except x=0 so yes ur on it, but like, do u just see this problem n knw, or do u go thru a guess n check process, what process do u go thru to get an answer for these types of problems?

OpenStudy (anonymous):

well the denominator can never be 0

OpenStudy (anonymous):

so x cannot be 0

OpenStudy (anonymous):

not much to guess at really

OpenStudy (anonymous):

just means if x were zero

OpenStudy (anonymous):

the value would be undefined

OpenStudy (anonymous):

hence a hole in your graph

OpenStudy (anonymous):

got it! ok i got one more...

OpenStudy (anonymous):

y=sqrt(2x+3)

OpenStudy (anonymous):

well since it is a square root is can never be negative

OpenStudy (anonymous):

so it would be \[(0,\infty)\]

OpenStudy (anonymous):

the graph of a square root function looks like

OpenStudy (anonymous):



OpenStudy (anonymous):

the books answer is, all X greater than or equal to -3/2

OpenStudy (anonymous):

yea thats right

OpenStudy (anonymous):

because

OpenStudy (anonymous):

y=sqrt(2x+3)

OpenStudy (anonymous):

to make the square root 0 x would have to be -3/2

OpenStudy (anonymous):

hence its still not a negative

OpenStudy (anonymous):

been up all night sorry

OpenStudy (anonymous):

getting tired

OpenStudy (anonymous):

but since -3/2 make it the sqrt of 0 thats the limit

OpenStudy (anonymous):

all other numbers have toi be greater then -3/2

OpenStudy (anonymous):

so if u plug in -3/2 "sqrt(2(-3/2)+3" this will equal 0?

OpenStudy (anonymous):

yes because the 2's cancel

OpenStudy (anonymous):

2x-3/2 = -3

OpenStudy (anonymous):

-3+3 = 0

OpenStudy (anonymous):

got it! so on these types of problems are u really just looking for w/e numbers will make it undefined?

OpenStudy (anonymous):

basically

OpenStudy (anonymous):

and if you cannot make it undefined

OpenStudy (anonymous):

its continuous from -inf to +inf

OpenStudy (anonymous):

ok well you have a good one, gotta get some rest now

OpenStudy (anonymous):

thanks!